OpenStudy (anonymous):

Pre-Cal Help!! Find domain of f: f(x)=(sqrt(2x-5))/(x^2-5x+4) The answer is [5/2, 4) U (4, infinity) I understand how to get the 5/2but where does the 4 come from? And It's been a yr and a half since I did college algebra, can someone pls re-explain how you get the bracket and parentheses? TYSM for any help!

hartnn (hartnn):

To find the domain, you need to EXCLUDE points which make DENOMINATOR =0 so, here, first you'll solve for \(x^2-5x+4=0\) which gives you \(x=4,1\) so, you have to exclude points 4 and 1 from the domain. note that, since you have already concluded \(x \ge 5/2\) (right?) excluding x=1 is unnecessary. so, you just need to exclude x=4 from the domain you got. from \(x \ge 5/2\) , domain\( =[5/2, \infty) \) but 4 cannot lie inside the domain of f(x) so we break this domain into 2 sub-domains and then take its union. First from 5/2 to 4, then from 4 to infinity so, final domain \(=[5/2, 4) U (4, \infty) \) now about brackets, [ and ] brackets are used when x can take that particular value (domain contains that value) here, 5/4 is included in domain, so there's a [ bracket in front of it. but 4 isn't included, so there are ( and ) brackets while writing the domain. does this make sense ?? ask if still any doubts.....

hartnn (hartnn):

Understood everyone clearly .?

hartnn (hartnn):

*everything

OpenStudy (anonymous):

why do we not exclude x=1?

OpenStudy (anonymous):

because denominator can't be 0 it doesn't exist in the defenition

hartnn (hartnn):

we did exclude x=1 when we say \(x \ge 5/2 ,\: x \ge 2.5 \implies x\ne 1 \) thats why we didn't have to exclude it again....

OpenStudy (anonymous):

That's right! I understand now! I really needed to refresh my mind on the algebra in here. Thank you so much. One more question...When will x ever be less than a value?

hartnn (hartnn):

For expressions like \(\sqrt{x-a}\), since term inside square root must be >= 0 \(x-a \ge 0 \implies x\ge a \) here we get x greater than a value. to get less than value, you just need an expression like \(\sqrt{a-x}\) and again since term inside square root must be >= 0 \(a-x \ge 0 \implies x\le a \)

OpenStudy (anonymous):

I see! Thank you for the explanations! You really helped.